OpenStudy (anonymous):

How do I know whether a system of linear equations has either 1 solution, many solutions, or no solutions?? without solving the system.

OpenStudy (amistre64):

look at the structure of the equations

OpenStudy (amistre64):

if the are multiples of each other they are the same line and all points are in common

OpenStudy (amistre64):

if they differ by a constant, they have no solutions since they are parallel lines that ride of into the sunset

OpenStudy (amistre64):

any other situation only has one solutions

OpenStudy (anonymous):

if you know about matrix determinants, then that would tell you, if the determinant is none zero, it has one solution, otherwise it has either no solution or many solutions

OpenStudy (amistre64):

im assuming its the run of the barrel, ax+by=c linears of course

OpenStudy (amistre64):

matrix sounds a bit to solvingish to me :)

OpenStudy (anonymous):

@amistre64, good point...

OpenStudy (anonymous):

well I'm in college remedial math. i'm very confused.. I learned y=mx+b .. for example.. i have these two equations, x=y-2, and y-x=3 ....How can I determine the solution by just looking at these two equations?

OpenStudy (anonymous):

I mean, how can i determine how many solutions it has.. or if it even has one

OpenStudy (amistre64):

youre asking how to do the math in your head without knowing how to do the math ....

OpenStudy (amistre64):

we know the constants are different, so it either has 1 or none

OpenStudy (amistre64):

if you can mentally put x and y to the same side; then the slope is the ratio of their coefficients such that -Xc/Yc = slope

OpenStudy (amistre64):

or did i trip on my own genuis lol

OpenStudy (amistre64):

x-y = -1/-1 = 1 -x+y = 1/1 = 1 .... parallels :) no solution